OpenStudy (anonymous):

The ellipse has a foci located at (3,1) and (7,1) and a major axis length of 10. What are the coordinates of the vertices?

OpenStudy (whpalmer4):

If the foci are at (3,1) and (7,1), that means we have a horizontal major axis, and the center is at ((7+3)/2,1) = (5,1). Our major axis length of 10 means that the vertices will be located along the line y = 1, and half the major axis length to the left and half the major axis length to the right. Therefore, our vertices are at (5-(10/2),1) = (0,1) and (5+(10/2),1) = (10,1).

OpenStudy (anonymous):

@whpalmer4 thank you sooo much!!!!! do you also happen to know how to find the length of the minor axis?